OpenStudy (anonymous):

y=(x-3)^2-25 to factored form

OpenStudy (alexandervonhumboldt2):

what to find?

OpenStudy (anonymous):

factored form :p

OpenStudy (anonymous):

First, I would suggest that you should expand \((x-3)^2\)

OpenStudy (alexandervonhumboldt2):

(x+2)*(x-8)

OpenStudy (alexandervonhumboldt2):

(x-3)^2-25=(x-3-5)*(x-3+5)=(x-8)*(x+2)

OpenStudy (alexandervonhumboldt2):

difference of squares a^2-b^2=(a-b)*(a+b)

OpenStudy (alexandervonhumboldt2):

can you fan me please?